Fermi National Laboratory


Accelerator Update for January 13, 2006 - January 18, 2006

Friday January 13

The day shift began with Operations monitoring store #4590, the Recycler Ring (RR) stash at 8.1 E10, Antiproton Source (Pbar) stack at 76.4 E10, and with NuMI taking beam. MiniBooNE is off for horn testing, MCenter is off for analyzing magnet repair.

Operations put the Recycler into shot setup at 9:46 AM to stash antiprotons.

Pbar returns to stacking at 10:48 AM.

At 11:02 AM, preparation begins for PBar and Booster access.

At 12:37 PM, Linac also goes into access.

Booster access is complete, beam returns to NuMI and MiniBooNe at 2:42 PM.

PBar access is complete, Pbar returns to stacking at 3:27 PM.

MCenter (MIPP) begins taking beam at 4:20 PM.

Meson begins experiencing control problems at 5:22 PM, beam is off to MIPP.

Store #4590 was terminated at 5:45 PM in order to replace a separator power supply.

At 5:47 PM, CDF and D0 both made accesses to their collision halls.

Both CDF and D0 accesses were completed by 9:48 PM.

Meson controls problems were resolved and beam returns to Meson at 11:13 PM.

Tevatron (Tev) studies begin at 11:28 PM.

Operations put the Recycler into shot setup at 11:45 PM to stash antiprotons.

Saturday January 14

The midnight shift began with the Tevatron in study mode, the Recycler Ring (RR) stash at 88.4 E10 and in shot setup, and with MiniBooNE and NuMI taking beam.

Pbar returns to stacking at 12:49 AM, beam also returns to MCenter.

Operations put the Recycler into shot setup at 5:28 AM to stash antiprotons.

Pbar returns to stacking at 6:24 AM, beam also returns to MCenter.

Tevatron shot setup begins at 7:33 AM.

The Tevatron quenched at 9:15 AM. Investigation and recovery are underway.

MiniBooNE turned off at 9:19 AM for a 6 hour access.

Beam to MIPP at 10:01 AM.

MiniBooNE is taking beam once again at 12:32 PM.

Operations put the Recycler into shot setup at 5:11 PM to stash antiprotons.

Pbar returns to stacking at 6:10 PM, MCenter is off due to Tevatron vacuum wook.

At 7:03 PM it was determined that the vacuum at A4 in the Tevatron is bad. This section of the Tev will need to be warmed up and the bad components will need to be replaced. Best estimate for Tevatron beam is 2 weeks.

Sunday January 15

The midnight shift began with the Tevatron being warmed up to repair vacuum leak in A4, the Recycler Ring (RR) stash at 129.8 E10, with the Pbar stack at 32.0 E10 and with MiniBooNE and NuMI taking beam.

Beam to MIPP at 12:53 AM.

Pbar begins stacking improvement studies at 11:09 AM.

Pbar returns to stacking at 8:28 PM.

Operations put the Recycler into shot setup at 10:58 PM to stash antiprotons.

Pbar returns to stacking at 11:44 PM, beam also returns to MCenter.

Monday January 16

The midnight shift began with the Tevatron being warmed up to repair vacuum leak in A4, the Recycler Ring (RR) stashing, with Pbar stacking, MiniBooNE, NuMI and MIPP taking beam.

Operations put the Recycler into shot setup at 3:12 AM to stash antiprotons.

Pbar returns to stacking at 4:12 AM, beam also returns to MCenter.

Pbar begins reverse proton studies at 8:39 AM.

At 10:55 AM, MIPP is being turned off for Transfer Hall access for Tevatron vacuum work.

At 2:30 PM , the Transfer Hall access is finished, MIPP is once again taking beam.

At 2:28 PM, D0 dropped their interlocks and went into supervised access.

Booster had frequent RF problems that caused beam downtime.

Tuesday January 17

The midnight shift began with the Tevatron A4 warmup, the Recycler Ring (RR) stash at 237.0 E10, and with MiniBooNE, NuMI and MIPP taking beam.

At 12:22 AM, Pbar studiers went home for the night.

Pbar studies begin at 7:50 AM.

MiniBooNE turns off at 8:26 AM.

MIPP turns off at 8:54 AM for analyzing magnet (MC7AN1) repair.

Booster was turned off at 10:09 AM to rack in BRF12, it was back on at 11:11 AM.

Tevatron A4 was warm at 3:06 PM. Leak testing/repair will begin.

MiniBooNE turns back on at 4:39 PM.

MIPP turns back on at 8:00 PM

Wednesday January 18

The midnight shift began with Tevatron A4 warm, the Recycler Ring (RR) stash at 218.5 E10, Pbar in study mode and with MiniBooNE, NuMI and MIPP taking beam.

Pbar studies were finished at 2:11 AM.

Pbar studies start up at 7:17 AM.

MIPP turns off at 7:24 AM for analyzing magnet (MC7AN1) repair.

The Plan for Wednesday

Pbar studies, keep the stash, run beam to NuMI and MiniBooNE.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 01/20/2006    email Fermilab